If you knew that the vertical intercept for a straight line was 15, that the slope was -.5, and that the independent variable had a value of 8, the value of the dependent variable would be:

Select one:

a. 8

b. 9

c. 10

d. 11
Answer #1

General Equation of the line is :-

y=mx+c

15 = C is the vertical intercept (Given)

-5 = m (which is the slope of the equation)

Thus the equation would be

y = -5x +15

y is the dependent variable and x is the independent variable.

thus if x=8

then, y= -5(8) + 15

=>y=-40 +15

= -25

This should be the answer, however its not given in the options.

Note:: Please check the either the question is wrong or the options are incorrect.